OpenStudy (anonymous):

Create your own 2-step equation with All Real Numbers as the solution. Show work and explain in words why this solution fits your equation.

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

So like 2x=10+2? Combine like terms to get 12 and then Divide by 2?

OpenStudy (anonymous):

I guess we'll never find out...

OpenStudy (anonymous):

I guess not.

OpenStudy (anonymous):

@rainbowtonic is what I said before an example of what you're looking for?

OpenStudy (anonymous):

Like this 4x+2 = 2 + 4x -2 -2 ------ ----------- 4x = 4x /4 /4 ------- -------- x = x So now you know that all real numbers will give you the correct answer to this equation. Because x=x no matter what number you put in for x.